Fermi National Laboratory
 

Accelerator Update for October 22, 2007 - October 23, 2007

 

Monday October 22

The day shift began with the Antiproton Source (Pbar) stacking (100E10), with Operations monitoring a Recycler stash (235.7E10), with the Main Injector (MI), Recycler, Pbar, and Tevatron (TeV) conducting studies, with the SciBooNE / MiniBooNE experiments taking beam, and with the Neutron Therapy Facility (NTF) treating patients.

At 8:27 AM, Operations reported that a TeV RF station (TRF7) was failing.

Operations reported at 9:46 AM that TeV experts brought the TeV down to 150 GeV and emptied it of all protons.

Operations reported at 9:51 AM that RF experts began repairs on TRF7.

At 10:06 AM, Operations reported that a MI RF station (MIRF01) had tripped. Experts investigated and had it back on 20 minutes later.

Operations halted the stacking at 10:06 AM and Pbar experts began reverse protons studies.

Operations reported at 10:07 AM that they had stashed the Recycler with 2 transfers of antiprotons.

TeV experts resumed their study at 10:57 AM.

At 11:18 AM, the TeV suffered a quench at TeV sector A1. The cause was a kicker (AAK2) prefire. Experts investigated.

At 12:10 PM, Operations prepared the TeV for an access. Cryo system experts reported that there was a stuck kautzky valve at TeV sector A13.

Operations prepared the Collider Detector at Fermilab (CDF) for access at 12:36 PM.

Operation prepared the Transfer Hall, sector A, for an hour access at 12:40 PM to repair the stuck kautzky valve.

NTF completed treating their last patient for the day at 1:03 PM.

Operations reported at 1:41 PM that the Recycler began transfers of antiprotons to MI.

At 6:37 PM, Operations reported a quench indication while turning things on. No heater firing units had fired. There was no quench. They also found a lead failure at TeV sector C4.

Pbar resumed stacking at 7:02 PM.

Operations put the TeV through a dry squeeze at 7:47 PM.

TeV studies resumed at 8:21 PM.

At 8:29 PM, the TeV suffered a quench at sector E3. An operator reported that beam had been injected on top of beam. TeV experts investigated.

Pbar experts began sending reverse protons to the Accumulator at 9:07 PM.

Operators reported at 11:30 PM that they had searched and secure Meson enclosures M03, 4, and 5.

Operations began tunes ups for transferring antiprotons from the Recycler to the TeV at 11:44 PM.

Tuesday October 23

The midnight shift began with Pbar conducting studies, with Operations monitoring the stash, with TeV experts preparing to accept antiprotons, and with Sci/MiniBooNE taking beam.

TeV experts completed a successful reverse proton tune up at 1:08 AM.

Operations reported at 1:42 AM that the first transfer of antiprotons was successfully injected into the TeV. Three more transfers had been made by 1:42 AM. Eventually, Operations established a 0x36 antiproton only store.

Operators charged the TeV abort kickers and aborted the TeV store at 2:44 AM.

A Recycler expert began a proton study at 2:44 AM.

Operations put the TeV through a dry squeeze at 3:51 AM.

Operations turned off beam to Sci/MiniBooNE at 5:03 AM.

Operations began preparing the TeV for access at 5:54 AM.

Operations began preparing Pbar for access at 5:55 AM.

Operations had the CDF and D-Zero (D0) collision halls ready for access by 6:01 AM.

Operations began preparing Booster for access at 6:04 AM.

Operations had MI prepared for access by 6:29 AM.

Operations had Linac prepared for access by 6:50 AM.

The Plans for Wednesday

The current status for today is as follows: Linac, Booster, Main Injector, Tevatron, Antiproton Source, CDF, and D0 are all in access. The D0 detector repairs are driving the shutdown.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/23/2007    email Fermilab